The sum of three numbers is 10 . Two times the smallest is 1 less than the largest, while the sum of the largest and smallest is 7 .
Answer by CubeyThePenguin(3113)

Let x < y < z

x + y + z = 10
2x = z - 1
x + z = 7

y = (x + y + z) - (x + z) = 10 - 7 = 3

2x - z = -1
x + z = 7

Add equations 2 and 3 to get 3x = 6, so x = 2 and z = 5.

(x, y, z) = (2, 3, 5)